By not negotiating the price of your first car purchase, you have a $1,000 "bad" money burn. How long would it take you to earn back the thousand dollars you just threw away if you take home $10.00 per hour??
thanks for the help!
A. 100 hours?
B. 10 hours?
C. 200 hours?
D. 20 hours?
It doesn't say how much hours I would work a week so that's why I am confused.

Answer by jim_thompson5910(35256)

(wage)*(number of hours) = total amount


($10 an hour)*(x hours) = $1000


10x = 100


10x/10 = 1000/10


x = 100


It will take you 100 hours (which is a little over 4 straight days of working nonstop)
------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------